# Tackling Long-Horizon Tasks with Model-based Offline Reinforcement Learning

### v1

@@ Appendix A Training Details
We report task-agnostic hyperparameters of our method in Table 7. We note that we use the same hyperparameters across all tasks, except $\tau$. We search the value of $\tau$ in $\{0.1,0.3,0.4,0.5\}$ and report the best value for the main experimental results. In addition, we report the exhaustive results in Tables 11 and 12, and summarize $\tau$ used in the main results in Table 8.
Table 7: Shared hyperparameters of LEQ.
@@
$\beta_{\text{EMA}}$ 1 Weight for critic EMA regularization
@@
Table 8: Task-specific hyperparameter $\tau$ of LEQ.
@@
We report the best hyperparameters of MOBILE for the AntMaze tasks in Tables 10 and 10. For MOBILE and MOBILE, we search the value of $c$ within $\{0.1,0.5,1.0,1.5\}$, as suggested in MOBILE [30], where $c$ is the coefficient of the penalized bellman operator:
@@ Appendix B Proof of the Policy Objective
$\hat{J}_{\lambda}(\theta)=\mathbb{E}_{\tau\sim p_{\psi},\pi_{\theta}}[(W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau))Q_{t}^{\lambda}(\tau)],$ (14)
@@
leads to optimizing a lower-bias estimator of $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
@@
To show this, we first prove that $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))]}$ is closer to $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$. For deriving the proof, we generalize this situation to have an arbitrary distribution $X$ and estimation $\hat{Y}$, which corresponds to $X=Q_{t}^{\lambda}(\tau),\hat{Y}=Q_{\phi}(\mathbf{s},\mathbf{a})$.
@@
Let $X$ be a distribution and $Y=E^{\tau}[X]$ be a lower expectile of $X$ (i.e. $0<\tau\leq 0.5$). Let $\hat{Y}$ be an arbitrary estimation of $Y$, and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$. If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimation of $Y$, then $|Y-\hat{Y}_{\text{new}}|\leq\frac{1-2\tau}{1-\tau}\cdot p(Y\leq X\leq\hat{Y})\cdot|Y-\hat{Y}|$.
@@
Without loss of generality, we assume $\hat{Y}\geq Y$. Then, we have $\hat{Y}_{\text{new}}\geq Y$. Thus,
$|\hat{Y}_{\text{new}}-Y|$
$=\hat{Y}_{\text{new}}-Y$
$=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{\qquad\qquad\qquad\qquad\qquad\qquad\;\;\;($\because$ Def. of $\hat{Y}_{\text{new}}$ and $Y$)}$
$=\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})X]}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{($\because$ Def. of $W^{\tau}(\cdot)$)}$
$=\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})X]-\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})])}$
$\leq\frac{1-2\tau}{\mathbb{E}[W^{\tau}(Y>X)]^{2}}\cdot(\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})X]-\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})]$
$=\frac{1-2\tau}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})X]-Yp(Y\leq X\leq\hat{Y}))$
$=\frac{(1-2\tau)p(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)$
$\leq\frac{(1-2\tau)p(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\hat{Y}-Y)$
$\leq\frac{1-2\tau}{1-\tau}\cdot p(Y\leq X\leq\hat{Y})\cdot|\hat{Y}-Y|$
Note that this theorem shows that the bias of the new estimation is always smaller than the original estimation, since $\frac{1-2\tau}{1-\tau}<1$ and $p(Y\leq X\leq\hat{Y})\leq 1$. If we plug in the distribution of $Q_{t}^{\lambda}(\tau)$ to $X$ and $\hat{Y}=Q_{\phi}(s_{t},\mathbf{a}_{t})$, then $Y=\mathbb{E}^{\tau}[X]=\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$, and we can show the desired result using the theorem: $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau)]}$ is closer to $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
@@
Here, the normalizing factor $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))]$ is non-differentiable with $\tau$. Specifically, the gradient is 0 everywhere (except $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}^{\lambda}(\tau)$). Thus, if we calculate the gradient of $\hat{Y}_{\text{new}}$, the gradient for the normalizing factor disappears. Therefore, we can omit the normalizing factor and get an equivalent formula $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]$ for gradient-based optimization.
Appendix C More Results
High variance in locomotion tasks.
When we train LEQ in locomotion tasks, we observe that our method often achieves $100$% success rates and then falls back to 00%, as shown in Figure 5. This is mainly because the learned models sometimes fail to capture failures (e.g. hopper and walker falling off) and predict an optimistic future (e.g. hopper and walker walking forward).
Figure 5: High variance during training. Our algorithm experiences oscillation on the performances due to optimistic imaginations near the initial states.
Results for all expectiles $\tau$.

### v2

@@ Appendix A Training Details
For state-based experiments, we report task-agnostic hyperparameters of our method in Table 9. We note that we use the same hyperparameters across all state-based tasks, except $\tau$. We search the value of $\tau$ in $\{0.1,0.3,0.4,0.5\}$ and report the best value for the main experimental results. In addition, we report the exhaustive results in Tables 15 and 16, and summarize $\tau$ used in the main results in Table 10.
For pixel-based experiments, we decrease the horizon length of DreamerV3 from $15$ to $5$, since we do not observe performance improvement with the longer imagination horizon (Appendix D), consistent with the finding from Lu et al. [2023]. Moreover, we remove the entropy bonus, as exploration is not required in offline RL. We search the value of $\tau$ in $\{0.1,0.3,0.4,0.5\}$ and report the best value in the main experimental results. All other hyperparameters follow the default settings of DreamerV3. We also report the exhaustive results in Table 17, and summarize $\tau$ in Table 13.
Table 9: Shared hyperparameters of LEQ in state-based experiments.
@@
$\omega_{\text{EMA}}$ 1 Weight for critic EMA regularization
@@
Table 10: Task-specific hyperparameter $\tau$ of LEQ in state-based experiments.
@@
We report the best hyperparameters of MOBILE for the AntMaze tasks in Tables 12 and 12. For MOBILE and MOBILE, we search the value of $c$ within $\{0.1,0.5,1.0,1.5\}$, as suggested in MOBILE [Sun et al., 2023], where $c$ is the coefficient of the penalized bellman operator:
@@
For MOPO in V-D4RL experiments, we search for $\lambda$ in $\{3,10\}$, as suggested in Lu et al. [2023], where $\lambda$ is the penalization coefficient in MOPO. Then, we report the best value in Table 14. For ROSMO, we use the hyperparameter specified in the official code.
@@
Table 13: Task-specific hyperparameter $\tau$ of LEQ in V-D4RL experiments.
Domain Task $\tau$
walker_walk random $0.5$
medium $0.3$
medium_replay $0.3$
medium_expert $0.1$
cheetah_run random $0.3$
medium $0.1$
medium_replay $0.1$
medium_expert $0.5$
Table 14: Task-specific hyperparameter $\lambda$ of MOPO in V-D4RL experiments.
Domain Task $\lambda$
walker_walk random $3.0$
medium $3.0$
medium_replay $3.0$
medium_expert $3.0$
cheetah_run random $3.0$
medium $3.0$
medium_replay $10.0$
medium_expert $10.0$
@@ Appendix B Proof of the Policy Objective
$\hat{J}_{\lambda}(\theta)=\mathbb{E}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}[(W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\mathcal{T}))Q_{t}^{\lambda}(\mathcal{T})],$ (13)
@@
leads to optimizing a lower-bias estimator of $\mathbb{E}^{\tau}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\mathcal{T})]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
@@
To show this, we first prove that $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))]}$ is closer to $\mathbb{E}^{\tau}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\mathcal{T})]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$ for most of the situations. For deriving the proof, we generalize the problem by considering an arbitrary distribution $X$ and its estimate $\hat{Y}$, which correspond to $X=Q_{t}^{\lambda}(\tau),\hat{Y}=Q_{\phi}(\mathbf{s},\mathbf{a})$. Then, we show $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ is closer to $Y$ than $\hat{Y}$. We split the case to two cases where $\hat{Y}\geq Y$ (Lemma 1) and $\hat{Y}\leq Y$ (Lemma 2) and show the effectiveness of $\hat{Y}_{\text{new}}$ instead of $\hat{Y}$ for each case.
Lemma 1.
Let $X$ be a distribution and $Y=E^{\tau}[X]$ be a lower expectile of $X$ (i.e. $0<\tau\leq 0.5$). Let $\hat{Y}$ be an arbitrary optimistic estimate of $Y$ (i.e., $\hat{Y}\geq Y$), and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$. If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimate of $Y$, then $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$.
Proof.
$|\hat{Y}_{\text{new}}-Y|$
$=\left\lvert\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\right\rvert\quad\text{\qquad\qquad\qquad\qquad\qquad\qquad\;\;\;($\because$ Def. of $\hat{Y}_{\text{new}}$ and $Y$)}$
$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\right\rvert\quad\text{($\because$ Def. of $W^{\tau}(\cdot)$)}$
$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]-\mathbb{E}[W^{\tau}(Y>X)\cdot X]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})])}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]-\mathbb{E}[W^{\tau}(Y>X)\cdot X]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})])}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]-Yp(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{(1-\tau)-(1-2\tau)\cdot p(Y\leq X)+(1-2\tau)\cdot p(Y\leq X\leq\hat{Y})}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{(1-\tau)-(1-2\tau)\cdot p(\hat{Y}\leq X)}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{\tau+(1-2\tau)\cdot p(X\leq\hat{Y})}\right\rvert$
$\leq\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{p(X\leq\hat{Y})}$
$\leq\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y$
$\leq\hat{Y}-Y=|\hat{Y}-Y|$
Lemma 1 shows that when the estimated value is optimistic ($\hat{Y}\geq Y$), the bias of the new estimate is always smaller than that of the original estimate. In the context of LEQ algorithm, the lemma tells that if the critic network ($\hat{Y}$) overestimates the lower-expectile of the target returns ($Y=\mathbb{E}^{\tau}[X]$), the surrogate loss ($Y_{\text{new}})$ compensates the overestimation of the critic values.
Unfortunately, when the estimated value is pessimistic ($\hat{Y}\leq Y$), there are some exceptional cases that the surrogate loss overcompensates the underestimation, resulting in an even larger error. The boundary for these cases is characterized in Lemma 2:
Lemma 2.
Let $X$ be a distribution and $Y=E^{\tau}[X]$ be a lower expectile of $X$ (i.e. $0<\tau\leq 0.5$). Let $\hat{Y}$ be an arbitrary conservative estimate of $Y$ (i.e., $\hat{Y}\leq Y$), and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$. If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimate of $Y$, then $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$, when $p(X\leq\hat{Y})\geq\frac{1}{2}(p(X\leq Y)-\frac{\tau}{1-2\tau})$.
Proof.
$|\hat{Y}_{\text{new}}-Y|$
$=\left\lvert\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\right\rvert\quad\text{\qquad\qquad\qquad\qquad\qquad\qquad\;\;($\because$ Def. of $\hat{Y}_{\text{new}}$ and $Y$)}$
$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)X]}{\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{($\because$ Def. of$W^{\tau}(\cdot)$)}\right\rvert$
$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)\cdot X]-\mathbb{E}[W^{\tau}(Y>X)\cdot X]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)])}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[\mathbbm{1}(\hat{Y}\leq X\leq Y)]-\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[\mathbbm{1}(\hat{Y}\leq X\leq Y)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)])}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{Yp(\hat{Y}\leq X\leq Y)-\mathbb{E}[\mathbbm{1}(\hat{Y}\leq X\leq Y)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{(1-\tau)-(1-2\tau)\cdot p(Y\leq X)-(1-2\tau)\cdot p(\hat{Y}\leq X\leq Y)}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{(1-\tau)-(1-2\tau)\cdot p(\hat{Y}\leq X)}\right\rvert$
$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{\tau+(1-2\tau)\cdot p(X\leq\hat{Y})}\right\rvert$
$=\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{\frac{\tau}{1-2\tau}+p(X\leq\hat{Y})}\right\rvert$
$=\frac{p(X\leq Y)-p(X\leq\hat{Y})}{\frac{\tau}{1-2\tau}+p(X\leq\hat{Y})}\cdot(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])$
$\leq Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X]$
$\leq Y-\hat{Y}=|\hat{Y}-Y|$
By combining Lemma 1, Lemma 2, we get Theorem 1 as below:
@@
Let $X$ be a distribution and $Y=E^{\tau}[X]$ be a lower expectile of $X$ (i.e. $0<\tau\leq 0.5$). Let $\hat{Y}$ be an arbitrary estimate of $Y$, and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$. If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimate of $Y$, then $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$ if the following condition holds:
$p(X\leq\hat{Y})\geq\frac{1}{2}(p(X\leq Y)-\frac{\tau}{1-2\tau}).$ (14)
@@
We combine the two cases dealt in Lemma 1, Lemma 2.
Here, $\hat{Y}\geq Y$ from Lemma 1 can be omitted, since the condition of Equation 14 includes the case of $\hat{Y}\geq Y$.
[…]
The condition of Equation 14 is a conservative bound applicable to any distribution, and $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$ holds across much broader regions in general. For example, Figure 8 shows that the inequality holds for 100%, 99.9% of the cases for normal, uniform distributions, respectively.
(a) Normal distribution
(b) Uniform distribution
Figure 8: Region of exceptions for the surrogate loss in realistic distributions. We visualize the cases where the surrogate loss is worse than the direct policy optimization, for two common distributions for $X$: (a) a normal distribution $\mathcal{N}(0,1)$ and (b) a uniform distribution $\mathcal{U}(0,1)$. We sample $\tau$ and $\hat{\tau}$ from arange($0,0.5,0.0025$) and compute $Y=\mathbb{E}^{\tau}[X]$, $\hat{Y}=\mathbb{E}^{\hat{\tau}}[X]$. The plots show the difference of errors $|\hat{Y}_{\text{new}}-Y|-|\hat{Y}-Y|$: positive values (red to yellow) indicate that the surrogate loss performs worse than the estimate, while negative values (green to blue) indicate that the surrogate loss outperforms the estimate. There is no exception for normal distribution, and only 39 out of 40000 cases are exceptional for uniform distribution.
@@
Here, we illustrate how we optimize $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]$ instead of $\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T})]}$. The normalizing factor $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))]$ is non-differentiable with $\mathcal{T}$ and the gradient is 0 everywhere (except $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}^{\lambda}(\mathcal{T})$). Thus, if we calculate the gradient of $\hat{Y}_{\text{new}}$, the gradient for the normalizing factor disappears. Therefore, we omit the normalizing factor and get an equivalent formula $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]$ for gradient-based optimization.
Appendix C Results for All Expectiles $\tau$